# The amplitude decay of a harmonic oscillator damped simultaneously by weak linear and nonlinear damping forces

### v1

@@ I Introduction
This paper is organized into five sections. In Section II, we present the theory and approximations we use and derive the first-order differential equation for the amplitude decay. In Section III, we obtain solutions in the case of damping by a combination of three damping forces. In Section IV, we obtain solutions in the case of damping by a combination of any two of the three considered damping forces. In Section V, we briefly comment on experimental setups in which our theoretical results could be tested and summarize important findings of the paper.
@@ II Basic theory, approximations we use, and equation of the amplitude decay
As an example of a damped harmonic oscillator, we consider a block of mass $m$ that oscillates back and forth along a horizontal surface under the influence of the restoring force of an ideal spring, i.e. $F_{res}(t)=-kx(t)$, where $k$ is the stiffness of the spring and $x(t)$ is the displacement of the block from the equilibrium position, and the damping force
@@
where $a(t)=d^{2}x(t)/dt^{2}$ is the acceleration of the block. The first term on the right-hand side of equation (3) models the force of sliding friction Lapidus ; AviAJP ; Grk2 ; AJPHinrich , while the second and third terms model, e.g., the influence of air resistance, since air resistance generally depends on terms proportional to velocity and square of velocity AJPpendulum ; AJPNelson ; Bacon ; Wang . In Section V we comment in more detail on physical systems where combinations of all three types of damping occur. The energy (potential plus kinetic) of the block-spring system is given by
@@
If we put $F_{d}(t)=0$ in (3) we get the equation of the undamped harmonic oscillator Resnick10 , with general solution $x(t)=A_{0}\cos(\omega_{0}t+\phi_{0})$, where $\omega_{0}=\sqrt{k/m}$ is the angular frequency of the undamped system, while $A_{0}$ and $\phi_{0}$ are constants (amplitude and initial phase) that are determined from initial conditions $\left(x(0)=x_{0},v(0)=v_{0}\right)$. The undamped system oscillates with conserved, i.e. constant, energy Resnick10 . For damped systems, i.e. for $F_{d}(t)\neq 0$, the energy is not conserved due to the power of the damping force $P_{d}(t)=F_{d}(t)v(t)$, and the energy dissipation rate is given by Young2020university
@@
All three types of damping are analyzed both theoretically and experimentally in Wang . Considering, e.g., the graphical representations of the experimental results presented in Wang , we can easily conclude that the defining characteristics of the weak damping regime, for each individual type of damping, are that the frequency and the initial phase remain approximately the same as in the undamped case, while the amplitude of the oscillations decreases very slowly over time, i.e. it decreases only slightly during one full oscillation (one period) of the system. Thus, in the case of weak damping with all three damping forces it is reasonable to assume that the solution of equation (3) is approximately of the form
@@
$v(t)=\frac{dx(t)}{dt}=-\omega_{0}A_{0}\left(f(t)\sin(\omega_{0}t+\varphi_{0})+\frac{\dot{f}(t)}{\omega_{0}}\cos(\omega_{0}t+\varphi_{0})\right)\,,$ (7)
where we used notation $\dot{f}(t)=df(t)/dt$ for simplicity. Since, by assumption, the amplitude changes very slowly compared to the oscillating part of (6), the rate of change of the function $f(t)$ over time has to be much slower than the rate of change of $\cos(\omega_{0}t+\varphi_{0})$, which is of the order $\omega_{0}$. We can conclude that $|\dot{f}(t)|\ll\omega_{0}$ holds for weak damping for all time, i.e., for $t\geq 0$ (we have put absolute value because we can expect the negative derivative of $f(t)$ since the amplitude decreases with time). Thus, due to $|\dot{f}(t)|/\omega_{0}\ll 1$, we can take that the velocity is approximately of the form
@@
as the final differential equation for the function $f(t)$. Since the function $f(t)$ monotonically decreases and the maximal value is $f(0)=1$, the weak damping condition $|\dot{f}(t)|\ll\omega_{0}$ $\forall t$ can be written as
@@
We note here that the condition (15) depends on the initial conditions due to $c_{0}\propto A_{0}^{-1}$ and $c_{2}\propto A_{0}$, i.e. in case of a damping force with terms non-linear in velocity, initial conditions play an important role in the overall dynamics of the system. The differential equation (13) can be solved by separation of variables and integration, i.e. we have to solve
@@ III Combinations of three damping forces
In Fig. 2(a) and (b) the solid blue curves show our approximate analytical solution (23) for the case $i=2$, i.e. the displacement $x_{2}(t)$, for two choices of coefficients $c_{n}$ and initial phases $\varphi_{0}=\{0,-\pi/2\}$. In Fig. 3(a) and (b) the solid blue curves show our approximate analytical solution (23) for the case $i=3$, i.e. the displacement $x_{3}(t)$, for two choices of coefficients $c_{n}$ and initial phase $\varphi_{0}=0$. In both Fig. 2 and Fig. 3, the dotted black curves show the corresponding numerical solutions. In Fig. 1- 3, we can see that our approximate analytical solutions give an excellent estimate of the duration of the motion, i.e. of the instant when the dynamics halts. The error of our approximate analytical solutions in the halting position compared to the corresponding numerical solutions can be at most $\Delta x=\pm\mu mg/k$, which amounts, e.g., to $\Delta x=\pm 1.6\%\,x_{0}$ for $\mu=0.01$ and the values of $m$ and $k$ that we use here, i.e. for $c_{0}=0.01\omega_{0}$.
@@
In Fig. 4 solid colored curves show our approximate energies (24), and black dotted curves show the corresponding numerically obtained energies. In Fig. 4(a) the chosen coefficients $c_{n}$ satisfy $c_{0}=c_{1}=c_{2}$, thus all solid colored curves in Fig. 4(a) correspond to energies $E_{1}(t)$ (i.e. (24) with $i=1$) calculated for five different values of coefficients $c_{n}$. In Fig. 4(b) the solid red curve corresponds to $E_{1}(t)$ with $c_{0}=0.05\omega_{0}$ and $c_{1}=c_{2}=0.01\omega_{0}$, the solid blue curve corresponds to $E_{2}(t)$ (i.e. energy (24) with $i=2$) with $c_{1}=0.05\omega_{0}$ and $c_{0}=c_{2}=0.01\omega_{0}$, and the solid green curve corresponds to $E_{1}(t)$ with $c_{2}=0.05\omega_{0}$ and $c_{0}=c_{1}=0.01\omega_{0}$. We see in Fig. 4(b) that the energy dissipation is the fastest in the case with dominant sliding friction.
@@ IV Combinations of two damping forces
V Discussion and conclusion

### v2

@@ I Introduction
This paper is organized into five sections. In Section II, we present the theory and approximations we use and derive the first-order differential equation for the amplitude decay. In Section III, we obtain solutions in the case of damping by a combination of three damping forces. In Section IV, we obtain solutions in the case of damping by a combination of any two of the three considered damping forces. In Section V, we analyze limitations of our approximate analytical solutions and comment on possible improvements to the presented approach. In Section VI, we briefly comment on experimental setups in which our theoretical results could be tested and summarize important findings of the paper.
@@ II Basic theory, approximations we use, and equation of the amplitude decay
Figure 1: Schematic representation of a block-spring system with a restoring force $F_{res}(t)$ and a (total) damping force $F_{d}(t)$. Here we show the time instant at which $x(t)>0$ and $v(t)<0$, i.e. at which $F_{res}(t)<0$ and $F_{d}(t)>0$. See text for details.
As an example of a damped harmonic oscillator, we consider the system schematically shown in Fig. 1, i.e. a block of mass $m$ that oscillates back and forth along a horizontal surface under the influence of the restoring force of an ideal spring $F_{res}(t)=-kx(t)$, where $k$ is the stiffness of the spring and $x(t)$ is the displacement of the block from the equilibrium position (set to $x=0$), and the damping force
@@
where $a(t)=d^{2}x(t)/dt^{2}$ is the acceleration of the block. The first term on the right-hand side of equation (3) models the force of sliding friction Lapidus ; AviAJP ; Grk2 ; AJPHinrich , while the second and third terms model, e.g., the influence of air resistance, since air resistance generally depends on terms proportional to velocity and square of velocity AJPpendulum ; AJPNelson ; Bacon ; Wang . In Section VI we comment in more detail on physical systems where combinations of all three types of damping occur. The energy (potential plus kinetic) of the block-spring system is given by
@@
If we put $\mu=b=D=0$ in (3), i.e. for $F_{d}(t)=0$, we get the equation of the undamped harmonic oscillator Resnick10 , with general solution
$x_{HO}(t)=A_{0}\cos(\omega_{0}t+\varphi_{0})\,,$ (5)
where $\omega_{0}=\sqrt{k/m}$ is the angular frequency of the undamped system, while $A_{0}$ and $\varphi_{0}$ are constants (amplitude and initial phase) that are determined from initial conditions $\left(x_{0},v_{0}\right)$, i.e. $A_{0}=\sqrt{x_{0}^{2}+(v_{0}/\omega_{0})^{2}}$ and $\varphi_{0}=\arctan\left(-v_{0}/(\omega_{0}x_{0})\right)$. The undamped system oscillates with conserved, i.e. constant, energy Resnick10 . For damped systems, i.e. for $F_{d}(t)\neq 0$, the energy is not conserved due to the power of the damping force $P_{d}(t)=F_{d}(t)v(t)$, and the energy dissipation rate is given by Young2020university
@@
All three types of damping are analyzed both theoretically and experimentally in Wang . Considering, e.g., the graphical representations of the experimental results presented in Wang , we can easily conclude that the defining characteristics of the weak damping regime, for each individual type of damping, are that the frequency and the initial phase remain approximately the same as in the undamped case, while the amplitude of the oscillations decreases very slowly over time, i.e. it decreases only slightly during one full oscillation (one period) of the system. Of course, such a conclusion can also be supported by purely theoretical considerations of the weak damping limit of exact analytical solutions in the case of viscous damping or sliding friction, while quadratic damping requires a more complex analytical or a simple numerical analysis, as we show in Section V. Thus, in the case of weak damping with all three damping forces, it is reasonable to assume that the solution of equation (3) is approximately of the form
@@
$v(t)=\frac{dx(t)}{dt}=-\omega_{0}A_{0}\left(f(t)\sin(\omega_{0}t+\varphi_{0})-\frac{df(t)}{dt}\omega_{0}^{-1}\cos(\omega_{0}t+\varphi_{0})\right)\,.$ (8)
Since, by assumption, the amplitude changes very slowly compared to the oscillating part of (7), the rate of change of the function $f(t)$ over time has to be much slower than the rate of change of $\cos(\omega_{0}t+\varphi_{0})$, which is of the order $\omega_{0}$. We can conclude that $|df(t)/dt|\ll\omega_{0}$ holds for weak damping for all time, i.e., for $t\geq 0$ (we have put absolute value because we can expect the negative derivative of $f(t)$ since the amplitude decreases with time). Thus, due to $|df(t)/dt|\,\omega_{0}^{-1}\ll 1$, we can take that the velocity is approximately of the form
@@
as the final differential equation for the function $f(t)$. Since the function $f(t)$ monotonically decreases and the maximal value is $f(0)=1$, the weak damping condition $|df(t)/dt|\ll\omega_{0}$ $\forall t$ can be written as
@@
The condition (16) depends on the initial conditions due to $c_{0}\propto A_{0}^{-1}$ and $c_{2}\propto A_{0}$, i.e. in case of a damping force with terms non-linear in velocity, initial conditions play an important role in the overall dynamics of the system. The differential equation (14) can be solved by separation of variables and integration, i.e. we have to solve
@@
We note here that an approximate solution of the form (7) passes through the equilibrium position at the same instants as the corresponding undamped solution (5), but it generally has turning points (i.e. maxima and minima) at earlier time instants compared to turning points of the undamped solution. This fact can be understood by qualitative consideration of the function (7), its derivative (8), and using the fact that $df(t)/dt<0$ for all $t\geq 0$. The function (7) describes oscillations that decay within the envelope given by ±$\pm A_{0}f(t)$. Let us assume that $A_{0}>0$ and that the function (7) has a turning point, e.g. a local maximum, at instant $t_{max}$, i.e. its derivative (8) is equal to zero at $t_{max}$. The maximum $x(t_{max})$ must be under the envelope, i.e. $x(t_{max})<A_{0}f(t_{max})$ must hold, because if it were on the envelope, since the envelope has some finite negative derivative at $t_{max}$ and function (7) has derivative equal to zero at that instant by assumption, immediately after $t_{max}$, i.e. for a slightly later instant, the function (7) would cross over the envelope due to the smaller magnitude of its derivative compared to the derivative of the envelope. Thus, the function (7) is under the envelope at $t_{max}$ and it will touch the envelope at some slightly later instant $t_{n}>t_{max}$ for which $x(t_{n})=A_{0}f(t_{n})<x(t_{max})$ and $\cos(\omega_{0}t_{n}+\varphi_{0})=1$ hold. We can conclude that $t_{n}$ is the instant of one of the turning points of the undamped solution (5) and that the corresponding turning point of the function (7) occurs at a slightly earlier instant. A similar argument can be applied to the turning points corresponding to the local minima of the function (7). Thus, the function (7), with $f(t)$ given by (17), generally (for any values of coefficients (15), and for any initial conditions) has turning points at earlier instants than the corresponding undamped solution (5), but for weak enough damping these differences will be negligible.
@@ III Combinations of three damping forces
In Fig. 3(a) and (b) the solid blue curves show our approximate analytical solution (24) for the case $i=2$, i.e. the displacement $x_{2}(t)$, for two choices of coefficients $c_{n}$ and initial phases $\varphi_{0}=\{0,-\pi/2\}$. In Fig. 4(a) and (b) the solid blue curves show our approximate analytical solution (24) for the case $i=3$, i.e. the displacement $x_{3}(t)$, for two choices of coefficients $c_{n}$ and initial phase $\varphi_{0}=0$. In both Fig. 3 and Fig. 4, the dotted black curves show the corresponding numerical solutions. In Fig. 2- 4, we can see that our approximate analytical solutions give an excellent estimate of the duration of the motion, i.e. of the instant when the dynamics halts. The error of our approximate analytical solutions in the halting position compared to the corresponding numerical solutions can be at most $\Delta x=\pm\mu mg/k$, which amounts, e.g., to $\Delta x=\pm 1.6\%\,A_{0}$ for $\mu=0.01$ and the values of $m$ and $k$ that we use here, i.e. for $c_{0}=0.01\omega_{0}$.
@@
In Fig. 5 solid colored curves show our approximate energies (25), and black dotted curves show the corresponding numerically obtained energies. In Fig. 5(a) the chosen coefficients $c_{n}$ satisfy $c_{0}=c_{1}=c_{2}$, thus all solid colored curves in Fig. 5(a) correspond to energies $E_{1}(t)$ (i.e. (25) with $i=1$) calculated for three different values of coefficients $c_{n}$. In Fig. 5(b) the solid red curve corresponds to $E_{1}(t)$ with $c_{0}=0.05\omega_{0}$ and $c_{1}=c_{2}=0.01\omega_{0}$, the solid blue curve corresponds to $E_{2}(t)$ (i.e. energy (25) with $i=2$) with $c_{1}=0.05\omega_{0}$ and $c_{0}=c_{2}=0.01\omega_{0}$, and the solid green curve corresponds to $E_{1}(t)$ with $c_{2}=0.05\omega_{0}$ and $c_{0}=c_{1}=0.01\omega_{0}$. We see in Fig. 5(b) that the energy dissipation is the fastest in the case with dominant sliding friction.
@@ IV Combinations of two damping forces
V Limitations of the presented approach and a comment on possible improvements
The presented approach is based on the assumption that the oscillation frequency and the initial phase in the weakly damped regime remain approximately the same as in the undamped case, while the amplitude gradually decreases, i.e. it is based on the assumption that a function of the form (7) can be taken as a good approximation for the solution of equation (3) in the case of weak damping. The results shown in the figures presented so far justify this assumption, i.e. the agreement between our approximate analytical solutions and numerical ones is better the weaker the total damping is (i.e. the smaller the sum $c_{0}+c_{1}+c_{2}$ is). So far we have not commented on the upper limit values of $c_{0}$, $c_{1}$, and $c_{2}$ for which our approximate solutions provide a good quantitative description of the dynamics of a harmonic oscillator damped simultaneously by all three damping forces. In this section, by considering the properties of solutions of equation (3) for each damping force individually, we provide an estimate on the limits of applicability of our approximate analytical solutions (24). Undergraduate students are typically familiar with the viscously damped harmonic oscillator, so we first consider the effect of viscous damping on the frequency and initial phase of the harmonic oscillations.
In case of viscous damping (i.e. for $b>0$, $\mu=D=0$), equation (3) has exact analytical solutions. We are interested here only in the underdamped, i.e. oscillatory, regime for which $c_{1}=b/(2m)<\omega_{0}$ holds. The corresponding exact solution for general initial conditions is given by (e.g. see Lelas2023 )
$x_{ex}(t)=\sqrt{x_{0}^{2}+\frac{(v_{0}+c_{1}x_{0})^{2}}{\omega^{2}}}e^{-c_{1}t}\cos(\omega t+\varphi)\,,$ (33)
where
$\omega=\omega_{0}\sqrt{1-\left(\frac{c_{1}}{\omega_{0}}\right)^{2}}$ (34)
is the damped angular frequency, and
$\varphi=\arctan\left(-\frac{v_{0}+c_{1}x_{0}}{\omega x_{0}}\right)$ (35)
is the initial phase. Thus, viscous damping reduces the frequency of oscillations compared to undamped oscillations, i.e. $\omega<\omega_{0}$, by an amount that depends only on $c_{1}$, while the initial phase $\varphi$ changes in relation to the initial phase of undamped oscillations $\varphi_{0}$ depending on both the initial conditions and the value of $c_{1}$. We can easily see that $\omega$ decreases slowly with increasing $c_{1}$, e.g. $\omega=0.995\omega_{0}$ for $c_{1}=0.1\omega_{0}$. The differences in the initial phases $\varphi$ and $\varphi_{0}$ in case $c_{1}>0$ are more subtle than the differences in $\omega$ and $\omega_{0}$. For example, for initial conditions $(x_{0}=0,v_{0}>0)$, we have $\varphi=\varphi_{0}=-\pi/2$ for any $c_{1}>0$, while for initial conditions $(x_{0}>0,v_{0}=0)$ and, e.g., $c_{1}=0.1\omega_{0}$ we have $\varphi=\arctan(-c_{1}/\omega)=-0.1$ and $\varphi_{0}=0$. Thus, the function $\arctan(-c_{1}/\omega)$ decreases faster than $\omega$ with increasing $c_{1}$. Therefore, in addition to changing the frequency, depending on the initial conditions, viscous damping can significantly change the initial phase, and both effects must be taken into account when estimating the range of values of $c_{1}$ for which (7) is a valid approximate form of the solution of equation (3). This issue has recently been addressed in LelasPezer where an approximate solution of the same form as (7) was used to model harmonic oscillations damped (only) with viscous damping, and it was shown that for $0<c_{1}\lesssim 0.1\omega_{0}$ the modeled solutions excellently describe the exact solution (33) with initial conditions $(x_{0}>0,v_{0}=0)$ and $(x_{0}=0,v_{0}>0)$. Thus, based on the analysis presented in LelasPezer , we can also take $c_{1}\approx 0.1\omega_{0}$ as the upper limit value of $c_{1}$ in our approximate solutions (24).
In case of sliding friction (i.e. for $\mu>0$, $b=D=0$) and initial conditions $(x_{0}>0,v_{0}=0)$, the exact analytical solution of equation (3) is (e.g. see Grk2 )
$x_{ex}(t)=\left(x_{0}-(2n-1)\frac{\mu mg}{k}\right)\cos(\omega_{0}t)-(-1)^{n}\frac{\mu mg}{k}\,$ (36)
where $n\geq 1$ is an integer that represents the number of half periods, i.e. for $20\leq t\leq T_{0}/$ 2 we take $n=1$, for $T_{0}/2\leq t\leq T_{0}$ we take $n=2$, etc., where $T_{0}=2\pi/\omega_{0}$. We have already commented on one shortcoming of our approach with respect to sliding friction, i.e. our approximate solutions (24) describe oscillations that halt at equilibrium position, while the exact dynamics with sliding friction generally does not halt at equilibrium position. Here we will comment on some other shortcomings of our solutions (24) with respect to sliding friction by analyzing solution (36). It is easy to show that the solution (36) has turning points (maxima and minima) at the instants $t_{n}=n\pi/\omega_{0}=nT_{0}/2$, i.e. at the same instants as the corresponding undamped solution $x_{HO}(t)=x_{0}\cos(\omega_{0}t)$. Thus, in terms of the time interval between successive turning points, the sliding friction does not change the frequency of the oscillations compared to the undamped oscillations. What is slightly harder to see is that the solution (36) passes through the equilibrium position somewhat later than the corresponding undamped solution and that this time difference increases in each subsequent half period. The corresponding undamped solution passes through the equilibrium position at instants $\tilde{t}_{n}=(2n-1)\pi/(2\omega_{0})$. If we put $t=\tilde{t}_{n}+\Delta t_{n}$ in (36) and equate it to zero, we get
$\Delta t_{n}=\frac{1}{\omega_{0}}\arcsin\left(\frac{1}{\frac{kx_{0}}{\mu mg}-(2n-1)}\right)\,.$ (37)
In case of weak damping, i.e. for $kx_{0}/(\mu mg)\gg 1$, in general we have some small value $\Delta t_{1}>0$ for $n=1$, and the values of $\Delta t_{n}$ increase in magnitude with increasing $n$. The argument of the the arcsine function in (37) must be less than or equal to one, so the maximum integer $n_{max}$ that can appear in (37) is the largest integer less than $kx_{0}/(2\mu mg)$. Thus, if the undamped solution passes through the equilibrium position at instants $\tilde{t}_{n}$, solution (36) will pass it at instants $\tilde{t}_{n}+\Delta t_{n}$. These ”time shifts” increase with $n$, and the largest is obtained for $n_{max}$. Accordingly, solution (36) passes through the equilibrium position $n_{max}$ times and, sometime after the last passage through the equilibrium position, halts at some finite displacement. As we commented in the last paragraph of Section II, approximate solutions of the form (7), unlike the solution (36), have turning points somewhat earlier than the undamped solution and pass through the equilibrium position at the same instants as the undamped solution. Given these differences and the aforementioned error of our approximate analytical solutions in the halting position, the sliding friction, i.e. the value of $c_{0}$, significantly affects the quantitative precision of our approximate description of the dynamics governed by equation (3). Still, in Fig. 2- 4 we can see that the quantitative agreement between our approximate analytical solutions and numerical solutions is fairly good for values $c_{0}\lesssim 0.03\omega_{0}$, and in Fig. 5 we see that our approximate energies (25) follow the behavior of the numerically obtained energies well even for $c_{0}=0.05\omega_{0}$. Furthermore, for $c_{0}\approx 0.03\omega_{0}$ (which corresponds to $\mu\approx 0.03$ for our choice of $m$, $k$, and $A_{0}$) the error of our solutions in the halting position can be at most $\Delta x\approx\pm 5\%\,A_{0}$. If we consider this value of $\Delta x$ to be acceptable, we can take $0<c_{0}\lesssim 0.03\omega_{0}$ to be the range of values of $c_{0}$ for which our solutions (24) describe quantitatively well the influence of sliding friction.
Figure 7: Numerical solutions of equation (3) with initial conditions $(x_{0}=A_{0},v_{0}=0)$ for $c_{2}/\omega_{0}=\{0.05,0.1,0.2,0.5\}$ and $c_{0}=c_{1}=0$. Vertical dashed lines are positioned at half periods of the corresponding undamped harmonic oscillator. See text for details.
In case of quadratic damping (i.e. for $D>0$, $\mu=b=0$), equation (3) is a nonlinear differential equation that cannot be solved analytically. The influence of quadratic damping on the frequency of harmonic oscillations can be studied by approximate analytical methods and by considering the time intervals between two consecutive turning points, i.e. half-cycles, but such an approach is mathematically demanding, e.g. see Smith ; AJPNelson . Here, for simplicity, we will use the numerical solutions of equation (3) to gain insight into the influence of quadratic damping on harmonic oscillations and estimate the upper limit value of $c_{2}$ for which our approach gives a good quantitative description of the dynamics. In Fig. 7 we show numerical solutions of equation (3) with initial conditions $(x_{0}=A_{0},v_{0}=0)$ for $c_{2}/\omega_{0}=\{0.05,0.1,0.2,0.5\}$ and $c_{0}=c_{1}=0$. We can see that for values $c_{2}/\omega_{0}=\{0.05,0.1\}$ the frequency of numerical solutions remains practically the same as in the undamped case, i.e. blue and red curves in Fig. 7 have all turning points positioned on the dashed vertical lines that denote instants of half periods of the undamped harmonic oscillator. Furthermore, we can notice that the numerical solution for $c_{2}=0.2\omega_{0}$, i.e. green curve, has a slight shift of the turning points to the right, while for $c_{2}=0.5\omega_{0}$, i.e. black curve, we can clearly see that the influence of quadratic damping is most pronounced during the first half cycle and that the time shift of the first turning point is ”transferred” to the subsequent turning points, i.e. after the first half cycle the numerical solution for $c_{2}=0.5\omega_{0}$ continues to oscillate with a frequency that is roughly the same as frequency of the undamped oscillator. Based on this brief analysis of the numerical solutions shown in Fig. 7, we can safely take $0<c_{2}\lesssim 0.1\omega_{0}$ to be the range of values of $c_{2}$ for which our solutions (24) provide a good quantitative description of the influence of quadratic damping on harmonic oscillations.
Figure 8: Solid blue curves show the solutions (24) with $A_{0}=0.2\,m$, $c_{1}=0.1\omega_{0}$ and $c_{2}=0.1\omega_{0}$ in all figures, while $c_{0}=0.05\omega_{0}$ in (a) and (b), and $c_{0}=0.03\omega_{0}$ in (c) and (d). Furthermore, $\varphi_{0}=0$, i.e. solutions with initial conditions $(x_{0}=A_{0},v_{0}=0)$ are shown in (a) and (c), while $\varphi_{0}=-\pi/2$, i.e. solutions with initial conditions $(x_{0}=0,v_{0}=\omega_{0}A_{0})$ are shown in (b) and (d). The dotted black curves show the corresponding numerical solutions.
In Fig. 8 we show results that suggest that we gave a reliable estimate of the upper bounds of the coefficients $c_{0}$, $c_{1}$, and $c_{2}$, i.e. our approximate solutions (24) provide a good quantitative description of the dynamics for $0<c_{0}\lesssim 0.03\omega_{0}$, $0<c_{1}\lesssim 0.1\omega_{0}$, and $0<c_{2}\lesssim 0.1\omega_{0}$. Once again, by comparing Fig. 8(a) and (b) with Fig. 8(c) and (d), we can see a strong influence of sliding friction on the precision of our approximate solutions, but even for our chosen upper limit value $c_{0}=0.03\omega_{0}$ in combination with upper limit values $c_{1}=c_{2}=0.1\omega_{0}$, i.e. for quite a high value $c_{0}+c_{1}+c_{2}=0.23\omega_{0}$, our approximate solutions describe the overall dynamics well and give a good estimate of the duration of motion.
Finally, we comment on possible improvements to the presented approach. As we have already mentioned, adding the viscous damping to the harmonic oscillator can affect the initial phase more than the frequency of the oscillations. Thus, the effect of viscous damping alone already suggests that more accurate approximate solutions could be obtained if, instead of a approximate solution of the form (7), we use
$x(t)=Af(t)\cos(\omega_{0}t+\Phi)\,,$ (38)
where $A\neq A_{0}$ and $\Phi\neq\varphi_{0}$, i.e. if we take into account that the initial phase changes, relative to $\varphi_{0}$, due to weak damping, and we keep the terms of first order in $|df(t)/dt|\omega_{0}^{-1}$ in velocity and energy, i.e. through the derivation of the function $f(t)$ by the time averaging of the energy dissipation rate. In such an approach, in addition to the initial conditions, $A$ and $\Phi$ will also depend on the damping constants $\mu$, $b$, and $D$, i.e. on the coefficients $c_{0}$, $c_{1}$, and $c_{2}$. Overall, these are nontrivial corrections to the approach used in this paper and are worth investigating on their own. The examples shown in Fig. 8(a) and (c) clearly indicate that taking into account the change in the initial phase could result in a better agreement between approximate analytical solutions and numerical solutions.
VI Discussion and conclusion